When a 100-m building is imploded, about how long does it take for a piece that breaks off from the roof to fall to the ground ?
Answer by fractalier(6550)

Well, the basic free fall equation (from rest, that is, and neglecting air resistance) is
y = (1/2)gt^2
Plug in what you know and solve for t...
100 = (1/2)(9.8)t^2
t^2 = 100 / 4.9 so that
t = 4.52 sec